OpenStudy (anonymous):

Which expression is equivalent to 3^2 ⋅ 3−5? fraction 1 over 3 to the power 3 fraction 1 over 3 to the power 7 fraction 1 over 3 to the power negative 3 fraction 1 over 3 to the power negative 7

OpenStudy (anonymous):

@haseeb96

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

Nnesha (nnesha):

and then change negative to positive exponent \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]